Entry
| Date
| Title
| Site
| Author
| #Graphics
|
163 | Sat 23-Oct-1999 | Cleaning Kryptons | stn3 | delany
|
160 | Fri 22-Oct-1999 | Cleaning Kryptons | | jirak
|
148 | Wed 20-Oct-1999 | Cleaned 5m KH2O at tower | tower |
|
143 | Tue 19-Oct-1999 | Change krypton gains from 10 to 1 in PAMS | | semmer
|
142 | Tue 19-Oct-1999 | Remounted KH2O sensor on Station 6 | stn6 | horst
|
126 | Sun 17-Oct-1999 | Krypton back on the air at 20 meters | tower | semmer
|
121 | Sun 17-Oct-1999 | Krpyton at 20m off the air at ~ 15:30 local time | tower | semmer
|
115 | Sun 17-Oct-1999 | h2o.k.20m back up | tower | semmer
|
114 | Sun 17-Oct-1999 | h2o.k.20m not working | tower | semmer
|
99 | Thu 14-Oct-1999 | Moved KH2O from Station 6 to 20m on 60m tower | stn6 | horst
|
79 | Mon 11-Oct-1999 | all stations have new krypton gain | | oncley
|
75 | Mon 11-Oct-1999 | Cleaned 5m KH2O | tower | horst
|
71 | Sun 10-Oct-1999 | cleaned stn 1-6 kryptons | stn4 | oncley
|
66 | Sat 09-Oct-1999 | looked at krypton voltage | stn1 | oncley
|
65 | Sat 09-Oct-1999 | cleaned krypton on stn3 | stn3 | oncley
|
57 | Fri 08-Oct-1999 | new krypton at 55m | tower | oncley
|
37 | Wed 29-Sep-1999 | Krypton spiking problem w/EVE A/D | stn3 | militzer
|
20 | Tue 21-Sep-1999 | krypton locations on 60 meter tower | | semmer
|
8 | Fri 03-Sep-1999 | How to Change PAM Krypton Calibration Files | | militzer
|
7 | Fri 03-Sep-1999 | Krypton Calibrations for PAM / Config.dat File | | militzer
|
5 | Thu 02-Sep-1999 | new kryptons tested | | oncley
|
- 163: Kryptons, Site stn3, Sat 23-Oct-1999 11:53:45 CDT, Cleaning Kryptons
Between 1100 and 1145, the Krypton hygrometers at stations 2, 5, 4 and 3
were cleaned using water wetted Q-tips and were then dried using Q-tips
- 160: Kryptons, Site , Fri 22-Oct-1999 11:51:50 CDT, Cleaning Kryptons
Cleaned Kryptons 1 and 6. Radiometer domes on station 1 were cleaned. Significant road dust was on the domes. Mouse also in the rain gauge.
- 148: Kryptons, Site tower, Wed 20-Oct-1999 14:56:16 CDT, Cleaned 5m KH2O at tower
I cleaned the windows of the KH2O at 5m at the main tower, using
a q-tip and distilled water. I intended to also clean those on the
PAM stations, but realized that the short ladder was not adequate
(duh!). Since I was too lazy to shlep the large ladder around to
all the PAM stations, I bagged the whole thing.
- 143: Kryptons, Site , Tue 19-Oct-1999 19:23:37 CDT, Change krypton gains from 10 to 1 in PAMS
Tom decided it was better to get some data with poor
resolution than no data at all! So we changed the
gain setting for the kryptons back to 1.
A new set of krypton.dat files were loaded into
each PAM station so if we want to change back to
a gain of 10 the files are ready to go.
file names: Kxxxxg1.dat - gain of 1
Kxxxxg10.dat - gain of 10
where "xxxx" is the sensor serial number.
NOTE: only the files for the sensor at the PAM
station have been downloaded.
This change was made from 1530 to 1615 CDT, Oct 19.
- 142: Kryptons, Site stn6, Tue 19-Oct-1999 19:11:56 CDT, Remounted KH2O sensor on Station 6
Replaced the KH2O sensor on Station 6. This had been mounted at 20m
on the 55m tower for several days. We lowered the mast around 1835 CDT
and raised it again around 1900 CDT, Oct 19.
This means that the h2o.k.20m on the tower no longer exist!
However, it is still in the config files for ops3.
- 126: Kryptons, Site tower, Sun 17-Oct-1999 19:19:37 CDT, Krypton back on the air at 20 meters
The krypton at 20 meters is back on the air. It is the
primary h2o sensor. The CO2 was also remounted and is
on the air. The boom was lowered at 18:25 CDT.
- 121: Kryptons, Site tower, Sun 17-Oct-1999 14:49:42 CDT, Krpyton at 20m off the air at ~ 15:30 local time
- 115: Kryptons, Site tower, Sun 17-Oct-1999 10:01:46 CDT, h2o.k.20m back up
The krypton at 20m is back on the air. The analog
signal cable was plugged into the wrong channel.
- 114: Kryptons, Site tower, Sun 17-Oct-1999 09:35:09 CDT, h2o.k.20m not working
The 20m krypton was not working throughout the night.
I may have incorrectly connected it back up after the
breakout box test.
- 99: Kryptons, Site stn6, Thu 14-Oct-1999 18:07:39 CDT, Moved KH2O from Station 6 to 20m on 60m tower
Removed the KH2O from Station 6 at 15:30 local time, Oct 14, and
installed at 20m because we moved the Station 6 sonic to 20m when
the latter died this moring. Calibrated data not available until
data system was reconfigured at 17:30. It was assumed from logbook
entry #19 that this was S/N 1393. From the calibration sheet in
the case, we used Kw = -0.145, Intercept = 4388, Path = 1.29.
- 79: Kryptons, Site , Mon 11-Oct-1999 18:51:22 CDT, all stations have new krypton gain
I've just loaded new krypton.dat files into all stns. This changes the
krypton voltage range to +/-1V, which hopefully will be enough, but also
increases our resolution by a factor of 10 to ~0.5mv.
This change also (finally) configured stn1 to have pp and not stn6.
We will have to change ops (during an IOP!) to accomodate this change
in the kryptons. I'll do that now.
- 75: Kryptons, Site tower, Mon 11-Oct-1999 11:08:28 CDT, Cleaned 5m KH2O
Cleaned 5m krypton hygrometer windows between 10:45 and 10:50 local time,
October 11.
- 71: Kryptons, Site stn4, Sun 10-Oct-1999 14:28:45 CDT, cleaned stn 1-6 kryptons
I cleaned all kryptons today, which brought up the signals significantly.
All were cleaned by lowering the masts. I disconnected the krypton signal
while the mast was down, which should indicate when I did this - approximately
10-1pm today. I worked on the stations in the order: 6, 4, 1, 3, 5, 2.
The sonics at 1 and 5 were loose on the booms, so I tightened the screws.
- 66: Kryptons, Site stn1, Sat 09-Oct-1999 21:36:42 CDT, looked at krypton voltage
I just disconnected the krypton from stn1 for a few minutes while checking
its voltage level (quite low: 0.090 V). It is reconnected now. This
done about 2115 local.
- 65: Kryptons, Site stn3, Sat 09-Oct-1999 16:01:51 CDT, cleaned krypton on stn3
I just reached up to clean the krypton on stn3 about 10 min. ago. It may
take a few minutes for this sensor to recover.
This done because Tom had noticed low fluxes on this sensor.
- 57: Kryptons, Site tower, Fri 08-Oct-1999 16:21:01 CDT, new krypton at 55m
By swapping things, we found that the head of S/N 1133 is responsible for
the 50s noise (and generally bad signal) from h2o.55m. We've swapped the
whole sensor with S/N 1389. We'll update cal_files soon.
P.S. Obviously, we (Tom and I) were up the tower -- from about 3:00-4:10.
- 37: Kryptons, Site stn3, Wed 29-Sep-1999 16:01:29 CDT, Krypton spiking problem w/EVE A/D
Krypton data from sensor 1391 was glitching as shown on the
cockpit display. It was the only sensor showing this trend.
A check running 'adflag' of the station revealed apparent
dropped a/d counts on the first cycle of the 20hz krypton task
after a reading taken of the .5hz power monitor parameters.
Approx 50 counts would be missed.
This is symptomatic of the a/d problem usually experienced
by the power monitor which can have a reduced vbatt reading.
Station 3 electronics box has been the poorest performing
PAM station in terms of responsiveness with the fastout and
high speed ingest processing. A couple of vx errors were
observed. This is probably a timing problem between the two
analog tasks and the device driver. It is likely either
the a/d board, dcom332 cpu or the backplane or a combination.
A/D board swap to station 1
From station 3 board 439220 removed and swapped into sta 1
From station 1 board 439176 removed and put in sta 3.
This should reveal whether the glitching problem follows the
A/D board. I don't think it will.
- 20: Kryptons, Site , Tue 21-Sep-1999 07:16:55 MDT, krypton locations on 60 meter tower
Serial number 1101 was installed at 5 meter height.
The calibration coefficients are:
path - 1.011cm
Kw - -0.137
V0 - 2677
Serial number 1133 was installed at 55 meter heigth.
The calibration coefficients are:
path - 1.434
Kw - -0.153
V0 - 2886
- 8: Kryptons, Site , Fri 03-Sep-1999 13:03:55 MDT, How to Change PAM Krypton Calibration Files
How To Change PAM Station Configuration when swapping Kryptons
between stations:
--------------------------------------------------------------
The steps basically amount to updating the "krypton.dat" file on
the PAM station by using the Rocky computer.
It is optional whether or not to also update the records on the base.
0) Physically swap in the new sensor.
1) Update the PAM station "krypton.dat" file
Use the rocky to interact with the station:
EVE> entersys
System # xmodem -yrt
'On-Rocky' send appropriate file such as 'Kryp1393.dat'
System EVE# copy kryp1393.dat krypton.dat
System EVE# start
And verify it works checking the id and data EVE sees.
Important Note: the Rocky is setup with the appropriate files to
send but you must select the proper version to use. This is
because the krypton files include a declaration for ingesting
the 20-hz delta-pressure sensor for stations 4,5 and 6.
For stations 4,5,6:
use the krypton file with deltaP included, select
c:\pam\cases99\krypton\with_dp\kryp13xx.dat
For stations 1,2,3:
use the krytpon file without deltaP included, select
c:\pam\cases99\krypton\no_dp\kryp13xx.dat
The krypton files for EVE include serial numbers 1389 thru 1395.
2) Update the Base Station Records (THIS IS OPTIONAL AND NOT REALLY NEEDED.
IT IS PREDICATED ON AN EARLIER METHOD OF CREATING PAM STATION CONFIGURATIONS
WHICH HAD THE KRYPTON DECLARATIONS IMBEDDED IN THE MAIN FORMAT ITSELF):
Edit file: $PAM/config/projects/CASES99
Change dated entry for the appropriate station and add a new dated entry
for the Krypton being swapped in something like this:
# STA_ID LAT LONG ELEVATION START_TIME STOP_TIME KRYPTON_SN LN_V0 1/PATHL_KW BIAS
1 37:40:00.00 -96:43:00.00 460 08/25/1999,00:00:00 10/05/1999,11:55:00 1389 8.518 -5.200857 0.0
1 37:40:00.00 -96:43:00.00 460 10/05/1999,12:00:00 12/01/1999,00:00:00 1393 8.387 -5.346163 0.0
#------------
Note: The date time must be formatted with 2-digit month/day/hour/min/sec
fields as shown
Note: You can always tweak the file time values later.
Create new station configuration files:
cd $PAM/projects/CASES99/sta_cfgs
evemake 73
Note: this assumes you are using configuration format #73. That may or
may not change during the project.
Note: evemake creates files called: s01_0073.dat.990825
for each station, format# and start date.
These files can alternately be copied to the Rocky and the
PAM station and renamed to 'config.dat', however the procedure
outlined above should be more straightforward.
- 7: Kryptons, Site , Fri 03-Sep-1999 11:43:05 MDT, Krypton Calibrations for PAM / Config.dat File
Krypton calibrations dated from Campbell 8/25/99 added to:
$PAM/config/projects/CASES99
For sensors 1389-1395.
PAM configuration files created assuming the following
station / sensor associations for now:
1 1389
2 1390
3 1391
4 1392
5 1393
6 1394
See Also: How to change PAM Krypton Cal Files
- 5: Kryptons, Site , Thu 02-Sep-1999 16:25:00 MDT, new kryptons tested
I just connected all of the new kryptons to cosmos:108 in serial number
(1389-1395) order. This was done between 15:50 - 16:25 local. All read
between 0.62 - 0.72 V, which was lower than what I expected, but still
usable. The first one, 1389, I tried cleaning with de-ionized water, but
actually reduced the output from 0.72 to 0.68 V. I cleaned it again and
got up to 0.74 V. Thus, I didn't try cleaning any of the others.